OpenStudy (anonymous):

write the equation in point slope form m=-4, (8,6)

OpenStudy (ddcamp):

Point slope form: y-y₁ = m(x-x₁) Where m is the slope and (x₁, y₁) is a point on the line.

OpenStudy (anonymous):

What is the answer

OpenStudy (jh3power):

y=-4x+38

OpenStudy (ddcamp):

(x₁, y₁) = (8,6); m=-4 Plug those values into the point slope equation: y-y₁ = m(x-x₁)

OpenStudy (ddcamp):

@jh3power That's slope intercept form.

OpenStudy (jh3power):

what's the difference...

terenzreignz (terenzreignz):

Difference is that the point slope form is nicer, but not the one that's specifically required in the question :D